OpenStudy (anonymous):

How to integrate 3 functions?

OpenStudy (anonymous):

Find \[\int\limits_{}^{} (-11e^{-t^{3}} * (\sin(x))^{4} d/dx) d/dx\]

OpenStudy (anonymous):

Oops It's -11e^x^3

OpenStudy (anonymous):

But I did find the derivative inside the integral which is... \[-44e^{-t^{3}} \sin^{3} x \cos(x)\]

OpenStudy (anonymous):

And then when you simplify, you get... \[-44 \int\limits_{}^{} e^{-x^{3}} \sin^{3} x \cos x dx\]

OpenStudy (anonymous):

^This is the step I'm stuck on...

ganeshie8 (ganeshie8):

Hey take a screenshot of actual problem and attach if psble

OpenStudy (anonymous):

Sure!

OpenStudy (irishboy123):

you are integrating a derivative here.

OpenStudy (anonymous):

Yes but when you see my work above ^ we end up with 3 functions inside the integral! :)

ganeshie8 (ganeshie8):

don't you see the integral and derivative kill each other out leaving you with the function inside so you don't have to do any work

ganeshie8 (ganeshie8):